### v1

# TabiBERT:A Large-Scale ModernBERT Foundation Model and Unified Benchmarking Framework for Turkish

Abstract
The introduction of BERT established encoder-only Transformer models as a foundational paradigm in natural language processing, motivating extensive research on architectural design and pre-training strategies. Subsequent work addressed limitations of early encoder models through improvements in computational efficiency, training stability, context length, and positional encoding. ModernBERT consolidated these improvements by integrating rotary positional embeddings, FlashAttention-based attention implementations, and refined normalization schemes to support longer contexts and more stable training. Unfortunately, the Turkish NLP landscape lacks a monolingual encoder trained from scratch that systematically incorporates these modern architectural advances.
This work introduces TabiBERT, a monolingual Turkish encoder based on the ModernBERT architecture that is trained from scratch on a large, carefully curated corpus. TabiBERT is pre-trained on one trillion tokens sampled from an 84.88B token multi-domain corpus comprising web text (73%), scientific publications (20%), source code (6%), and mathematical content (0.3%). The model supports a context length of 8,192 tokens, which is 16x of the original BERT models. It achieves an inference speedup of up to 2.65 times while reducing GPU memory consumption – allowing for larger batch sizes during training and inference. To support rigorous and reproducible evaluation, we introduce the TabiBench benchmark with 28 datasets across eight task categories with standardized splits and evaluation protocols. Performance is summarized using a GLUE-style macro-average score. On TabiBench, TabiBERT attains a score of 77.58, outperforming BERTurk by 1.62 points and establishing new state-of-the-art results on five of eight task categories, with particularly strong gains on question answering (+9.55 points), code retrieval (+2.41 points), and document retrieval (+0.60 points). When compared with task-specific prior best results, including specialized models such as TurkishBERTweet, TabiBERT achieves an average improvement of 1.47 points, indicating robust cross-domain generalization. We release model weights, training configurations, and evaluation code to provide a transparent and reproducible foundation for future Turkish encoder research.
1 Introduction
The emergence of pre-trained transformer encoder models has become central to contemporary natural language processing, providing general-purpose representations that support a wide range of downstream tasks. The architectural framework introduced by Bidirectional Encoder Representations from Transformers (BERT) (Devlin2019BERT) has been a foundational component of natural language understanding, enabling the induction of deeply contextualized linguistic features through self-supervised learning on large-scale corpora. BERT was first trained on large English text corpora and achieved strong baseline performance on a wide range of natural language processing tasks, serving as the basis for many later pre-trained language models.
The success of BERT motivated, and thus was followed by extensive research, leading to resource-efficient variants (Liu2019RoBERTa; Sanh2019DistilBERT; Jiang2020ConvBERT; Lan2020ALBERT; Clark2020ELECTRA), domain-specialized models (Beltagy2019SciBERT; Huang2019ClinicalBERT; Lee2020BioBERT; Chalkidis2020LegalBERT), and multilingual architectures designed to support cross-lingual transfer (Pires2019MultilingualBERT; Vulic2020ContinualLearning; Conneau2020XLM-R). Initial refinements focused on pre-training procedures that improved performance without changing the core architecture. For example, RoBERTa (Liu2019RoBERTa) achieved improved performance via larger and more diverse pre-training corpora, dynamic masking strategies, optimized training regimes, and by eliminating the next sentence prediction objective without altering the original transformer architecture. Several works concurrently focused on efficiency through knowledge distillation (Sanh2019DistilBERT), parameter sharing strategies (Lan2020ALBERT), and alternative pre-training objectives (Clark2020ELECTRA), enabling computational gains while maintaining or improving performance.
Subsequent architectural innovations sought to address BERT’s quadratic complexity, which stems from the self-attention mechanism and limits sequence length and computational efficiency. Beltagy2020Longformer introduced sparse attention to enable longer sequences, while IO-aware implementations significantly reduced memory overhead (Dao2022FlashAttention). Positional encoding improvements such as Rotary Position Embeddings (RoPE) (Su2024Roformer) resulted in improved length extrapolation compared to BERT’s original absolute positional embeddings. Normalization strategies were introduced to improve training stability for deeper architectures (Xiong2020LayerNorm; Zhang2019RootMeanSquare). These accumulated improvements were synthesized into ModernBERT, an updated encoder-only architecture that incorporates rotary positional embeddings, FlashAttention technology, alternating attention patterns, and refined pre-training procedures (Warner2024ModernBERT). Its architecture supports context windows up to $8,192$ tokens, sixteen times BERT’s original context length of 512, with improved computational efficiency that addresses limitations stemming from fine-tuning instability and restricted sequence length (Xiong2020LayerNorm; Liu2019RoBERTa).
While these architectural advances have been widely adopted for high-resource languages, Turkish NLP remains primarily centered on earlier architectural paradigms. BERTurk (Schweter2020BERTurk) laid the foundation for transformer-based Turkish NLP as the first monolingual BERT model pre-trained from scratch and publicly released. Its widespread adoption across diverse downstream tasks has made it the de facto standard for Turkish language understanding. Following, subsequent models such as ConvBERTurk and various ELECTRA adaptations have emerged on community platforms, although they frequently prioritized immediate availability over technical documentation. Critical specifications regarding corpus composition, preprocessing pipelines, and systematic evaluation protocols are often underspecified, which impedes reproducibility and limits the ability of the research community to build systematically on prior work.
Recent large-scale multilingual efforts have incorporated Turkish alongside other languages. Multilingual BERT (mBERT) (Devlin2019BERT) and XLM-R (Conneau2020XLM-R) share parameters and vocabulary across typologically diverse languages, which dilutes representational capacity for individual languages (Conneau2020XLM-R; Vulic2020ContinualLearning). For morphologically rich languages, multilingual tokenizers produce inefficient segmentations, leading to increased token fertility, longer effective sequences, and reduced vocabulary utilization (Rust2021Tokenizer; Hofmann2022Tokenization).
Table 1: Turkish BERT-based models across scale, training data, architectural features, and benchmark performance.
Model # Params (M) Data Size (GB) Context Length Code & Math Flash Attn. TabiBench Score
TurkishBERTweet $163$ $110$ $512$ $67.48$
YTU-Cosmos-BERT $111$ $75$ $512$ $72.26$
BERTurk $110$ $35$ $512$ $75.96$
TabiBERT $149$ $426$ $8,192$ $77.58$
This work identifies and addresses two distinct gaps in Turkish NLP research. First, an architectural gap exists: no monolingual Turkish encoder has been trained from scratch to incorporate the efficiency, stability, and representational improvements synthesized in ModernBERT. Second, an evaluation gap persists due to the lack of a standardized, reproducible benchmarking framework, hindering systematic comparison across model generations and architectural paradigms.
We address these gaps by introducing TabiBERT, a monolingual Turkish encoder based on the ModernBERT architecture, and TabiBench, a unified evaluation benchmark with standardized protocols. TabiBERT is a monolingual Turkish encoder based on the ModernBERT architecture, trained from scratch on a large, carefully curated corpus. Across TabiBench, TabiBERT achieves an average score of 77.58 compared to the next best model BERTurk’s 75.96, representing an improvement of 1.62 absolute points (on a 0-100 scale, following GLUE-style macro-averaging) and establishing state-of-the-art results among Turkish-specific encoders on five of eight task categories. Performance gains are particularly substantial on retrieval tasks, with improvements of 9.55 points on question answering, 2.41 points on code retrieval, as well as 0.60 points on document retrieval. When compared against the previous best result for each task individually, TabiBERT achieves an average improvement of 1.47 points, demonstrating strong cross-domain generalization with a single general-purpose architecture. Beyond accuracy improvements, TabiBERT provides 2.65× faster inference and supports a context length of 8,192 tokens, sixteen times that of prior Turkish BERT models, while reducing GPU memory consumption to enable larger batch sizes during training and inference.
More recently, mmBERT (Marone2025mmBERT), a 307307M-parameter multilingual model with a vocabulary of $256,000$ tokens trained from scratch at large scale, demonstrates strong performance on Turkish tasks, achieving 79.26 on TabiBench (see Appendix A). However, this performance comes at the cost of significantly increased computational requirements and the need for large-scale multilingual training data. Moreover, mmBERT’s multilingual tokenizer produces substantially higher token fertility for Turkish texts, +41% compared to TabiBERT’s monolingual tokenizer, resulting in longer effective sequences and reduced computational efficiency for Turkish-specific applications (detailed in Section 5). We include mmBERT as a reference point representing a different model class with substantially different resource requirements and design objectives, while our work focuses on developing an efficient, monolingual encoder that incorporates modern architectural advances at a more modest scale suitable for research and practical deployment.
The main contributions of this work are as follows:
  • TabiBERT (TabiBERT-HF), a state-of-the-art Turkish encoder that bridges the architectural gap in the Turkish NLP ecosystem by adapting the ModernBERT architecture. The model was pre-trained for a total of 1 trillion tokens on a multi-domain Turkish corpus comprising 84.8884.88B tokens, including web text (73%), scientific publications (20%), source code (6%), and mathematical content (0.3%). TabiBERT features an $8,192$-token context length, representing a 15× increase over previous Turkish BERT-style models, and achieves up to 2.65× inference speedup. The model achieves state-of-the-art results among Turkish-specific encoders on five of eight benchmark task categories, with an average improvement of 1.62 absolute points over BERTurk and particularly strong gains on question answering (9.55 points) and retrieval tasks (2.41 points on code retrieval, 0.60 points on document retrieval), while significantly reducing memory overhead.
  • TabiBench (TabiBench), a unified and comprehensive evaluation framework designed to standardize Turkish NLU research by providing a reproducible pipeline including standardized data splits, preprocessing scripts, and evaluation configurations. TabiBench comprises 28 datasets across 8 task categories, including specialized domains such as code retrieval and academic text understanding that have not been systematically evaluated in prior Turkish NLP benchmarks.
  • The open-source release of the TabiBERT model weights and the TabiBench evaluation codebase (TabiBERT) to provide a transparent foundation for future research in Turkish language models and downstream applications. Detailed corpus composition statistics and preprocessing methodology are provided in Section 3.
The remainder of the paper is structured as follows: Section 2 reviews the relevant literature on encoder architectures and Turkish language models; Section 3 presents the corpus and outlines the data preparation procedures; Section 4 details the pre-training methodology and architectural choices; Section 5 reports a comprehensive evaluation of TabiBERT across the TabiBench benchmark; and Section 6 summarizes the findings, discusses limitations, and identifies directions for future research.
2 Related Work
The introduction of Bidirectional Encoder Representations from Transformers (BERT) (Devlin2019BERT) was a significant advancement in encoder-only transformer models. These models were trained on large English corpora and established strong performance baselines across diverse natural language processing tasks. They provided a pre-trained model that could be fine-tuned for various downstream tasks.
2.1 Encoder-only Transformer Models & Architectural Evolution
Following BERT, encoder-only transformer models were improved by various architectural innovations or training techniques, including changes to pre-training objectives, optimization strategies, and large-scale data and model scaling, leading to consistent performance gains across a wide range of downstream NLP tasks. RoBERTa (Liu2019RoBERTa) was trained with the same masked language modeling objective but with a larger training corpus and bigger model sizes, achieving state-of-the-art performance on various NLP tasks. Following, ALBERT (Lan2020ALBERT) introduced a number of mechanisms such as parameter sharing and a factorized embedding parameterization, resulting in a significant reduction in model size and memory consumption. This enabled training deeper, wider encoder models while still performing competitively on downstream benchmarks. ELECTRA (Clark2020ELECTRA) proposed a more sample-efficient pre-training approach based on replaced token detection, where a discriminator model learns to distinguish real input tokens from plausible alternatives generated by a small generator network, resulting in strong performance with substantially reduced pre-training compute. DeBERTa (he2021debertav1) introduced disentangled attention to decouple content and positional representations, improving modeling of relative position information in encoder-only transformers. The success of DeBERTa was followed by DeBERTaV2 (microsoft_debertav2_repo), which refined this approach by using improved mask decoding and pre-training strategies. Moreover, DeBERTaV3 (He2023DeBERTaV3) further enhanced training efficiency by integrating a replaced-token detection objective inspired by ELECTRA.
Subsequent encoder-only transformer works after DeBERTaV3 focused on improving pre-training efficiency by revisiting training objectives and modernizing the BERT architecture rather than attempting to introduce novel attention mechanisms. MosaicBERT (Portes2023Mosaicbert) worked on engineering a training recipe by combining architectural and systems optimizations such as Flash Attention (Dao2022FlashAttention), ALiBi positional biases (Press2021Train), Gated Linear Units (Shazeer2020Glu), dynamic unpadding (Zeng2022Boosting), and low-precision normalization. This engineering effort demonstrated that BERT pre-training can run significantly faster and more cost-effectively, while maintaining competitive downstream performance. In parallel, task-oriented pre-training strategies were explored, exemplified by RetroMAE (Xiao2022RetroMAE), which redesigned masked autoencoding to better support dense retrieval tasks using a BERT-style encoder. FlexiBERT (Tuli2023FlexiBERT) used neural architecture search to identify performant encoder configurations under different computational constraints, aiming to investigate architectural flexibility and efficiency trade-offs. Renewed interest in encoder-only models has led to modernized successors to BERT. ModernBERT (Warner2024ModernBERT) revisited the original BERT model, extending context length and updating architectural elements and design choices. NeoBERT (Breton2025NeoBERT) and OptiBERT (Dervishi2025TrainingOptiBERT) emphasized compute-optimal scaling and practical deployment. Together, these studies show a trend toward improving the efficiency and practicality of BERT-style encoder models, while maintaining strong performance on language understanding tasks.
2.2 Monolingual Turkish BERT Models
The success of transformer-encoder models motivated a series of BERT models known as BERTurk to be trained on Turkish data (Schweter2020BERTurk). Several variations of BERT, such as DistilBERT (Sanh2019DistilBERT), ConvBERTurk (Schweter2020BERTurkV2), and electra (Clark2020ELECTRA) were also trained on Turkish data by the same authors (Schweter2020BERTurk). These models were generally trained on a 3535GB corpus consisting of 4.44.4B tokens drawn from the Turkish oscar corpus (Abadji2022OSCAR), a Wikipedia dump [^1] An unspecified Wikipedia dump, presumed to date from approximately 2020., and various OPUS corpora (Tiedemann2012OPUS). Some models, like ConvBERTurk and electra, were also trained on the Turkish portion of the mC4 corpus (xue2021mt5massivelymultilingualpretrained; Schweter2020BERTurk) which is a cleaned version of the public web crawl data of Common Crawl. These models have been evaluated on various downstream tasks (such as part-of-speech tagging, named entity recognition, and question answering) where they generally outperform their multilingual counterparts (Schweter2020BERTurk; wu2019betobentzbecassurprising; Conneau2020XLM-R). Despite their popularity, these models were limited in transparency as they were not products of academic publications.
Following, two Turkish BERT models were introduced with academic publications. YTU-Cosmos-BERT (Kesgin2023YTU) utilized an uncased Turkish tokenizer and was trained on a corpus of an unknown number of tokens reported to be around 7575GB in size, that consists of text collected from the web, Turkish Wikipedia and novels. TurkishBERTweet (Najafi2023TurkishBERTweet) was tailored for social media. It also employed an uncased tokenizer, specifically designed for social media texts, by the addition of special tokens that can encode special information in tweets like hashtags, mentions, and URLs. It was pre-trained with a corpus of around 110110GB in uncompressed form, containing about 900 million tweets of 10 tokens on average, which totals to a sum of 99B tokens when extrapolated. Recently, a RoBERTa-based model for Turkish was published (ScheibleSchmitt2025SindBERT), which has been an improvement over the previous RoBERTa-based models for Turkish (TasRoberTURK2024) as it was trained on a larger corpus, using a larger batch size. However, it did not achieve a significant improvement over the previous models.
2.3 Multilingual BERT Models
The development of multilingual pre-trained language models began with Multilingual BERT (mBERT) (Devlin2019BERT), which extended the original BERT architecture to 104 languages by pre-training on a corpus obtained by concatenation of Wikipedia corpora. Despite lacking explicit cross-lingual objectives, mBERT demonstrated surprising zero-shot cross-lingual transfer capabilities (Pires2019MultilingualBERT). XLM (Conneau2019Cross) worked on enhancing cross-lingual alignment and introduced Translation Language Modeling (TLM) that leverages parallel corpora. XLM-RoBERTa (Conneau2020XLM-R) further scaled multilingual pre-training to 2.52.5TB of CommonCrawl data across 100 languages, achieving state-of-the-art performance using only masked language modeling without parallel data. Later studies proposed different pre-training approaches. InfoXLM (Chi2021InfoXLM) applied cross-lingual contrastive learning, ERNIE-M (Ouyang2021ERNIE) used back-translation masked language modeling to create pseudo-parallel data, and Unicoder (Huang2019Unicoder) combined several cross-lingual pre-training objectives. Architectural improvements have also been explored in multilingual encoder models. RemBERT (Chung2020Rethinking) decoupled input and output embeddings in order to better use the model capacity. mDeBERTa (He2023DeBERTaV3) extended DeBERTa’s disentangled attention to the multilingual setting and combined it with ELECTRA-style pre-training, resulting in consistent improvements over XLM-RoBERTa on cross-lingual benchmarks. More recently, XLM-V (Liang2023XLMV) addressed the vocabulary bottleneck by improving token allocation across languages with varying lexical overlap.
More recent work has further expanded the landscape of multilingual encoders. EuroBERT (Boizard2025EuroBERT) introduced multilingual encoders covering European and several widely spoken global languages. Similarly, mmBERT (Marone2025mmBERT), building on ModernBERT, incorporated long-context attention and efficient training strategies. It demonstrated strong performance on classification, embedding, and retrieval tasks, particularly for low-resource languages, while largely preserving the inference efficiency of ModernBERT despite its large vocabulary of $256,000$ tokens.
However, these broad multilingual approaches often struggle with the linguistic specificity required for morphologically complex languages, motivating the development of language-family-specific or language-specific models.
2.3.1 Curse of Multilinguality
A key challenge in multilingual modeling is the curse of multilinguality, a phenomenon where adding more languages to a fixed-capacity model eventually degrades per-language performance (Conneau2020XLM-R). It is demonstrated (Wang2020Negative) that this stems from negative interference between languages competing for shared model parameters, affecting both high-resource and low-resource languages. In a comprehensive study conducted over 10,000 models and across 250 languages (Chang2024Multilinguality), it is observed that while moderate multilinguality can benefit low-resource languages, particularly when linguistically similar languages are added, performance degrades for all languages as the number of languages increases, suggesting that massively multilingual pre-training may not be optimal for any language involved.
Several architectural approaches have been proposed to address this issue, such as modular designs that use language-specific adapters (Pfeiffer2022Lifting) and expert models trained on subsets of languages (Blevins2024Breaking). These findings motivate the development of more targeted models that focus on either monolingual models or models that focus on related language families rather than attempting to cover all languages in a single model.
2.4 ModernBERT
ModernBERT is one of the recent and most significant bidirectional encoder-only Transformer models. It upgrades the original BERT architecture by utilizing contemporary enhancements to support longer context lengths and improve efficiency and performance in natural language understanding tasks. It was pre-trained on a massive corpus of  2 trillion tokens of English text and code, enabling stronger generalization and suitability for both language and code understanding. ModernBERT introduces several architectural techniques aimed at improving efficiency, including Rotary Positional Embeddings (RoPE) Su2024Roformer, Flash Attention (Dao2022FlashAttention), unpadding (Zeng2022Boosting), GeGLU activations (Shazeer2020Glu), and alternating attention (Team2024Gemma). Together, these design choices reduce memory usage and inference time compared to conventional encoder models, while preserving strong performance on downstream tasks. These features make ModernBERT effective for classification, retrieval, semantic search, and other NLU tasks, and it serves as an efficient alternative to older BERT variants in research and production settings.
2.4.1 Language-Specific Variants
The success of ModernBERT has inspired several language-specific adaptations, demonstrating the effectiveness of specialized training over multilingual approaches. For example, ModernGBERT is a family of German encoder models (134M, 1B parameters) trained from scratch, incorporating ModernBERT’s architectural innovations (Ehrmanntraut2025ModernGBERT). Their work showed that dedicated encoders outperform multilingual models adapted via LLM2Vec (Behnamghader2024Llm2vec) in terms of both performance and parameter-efficiency, validating the approach of language-specific model development.
llm-jp-modernbert is a similar ModernBERT-based model trained on a large-scale Japanese corpus with $8,192$ token context length (Sugiura2025ModernBERTJapanese). While achieving competitive results on fill-mask evaluations, the model highlighted the importance of high-quality, language-specific pre-training data for optimal performance in downstream tasks.
2.4.2 Domain-Specific Applications
The ModernBERT architecture has proven particularly effective for specialized domains that require robust long-context understanding and efficient representation learning. Clinical ModernBERT was developed as a transformer-based encoder pre-trained on biomedical literature, clinical notes, and medical ontologies, incorporating large-scale resources such as PubMed abstracts and MIMIC-IV clinical data (Lee2025ClinicalModernBERT). Building on ModernBERT’s architectural upgrades, most notably long-context support and efficient attention mechanisms, this domain-adapted model demonstrated strong performance in producing semantically rich representations for long-document biomedical tasks, including clinical classification, retrieval, and entity recognition.
BioClinical ModernBERT further extended this line of work through continual pre-training on the largest biomedical and clinical corpus to date, comprising more than 53.5 billion tokens (Sounack2025BioClinicalModernBERT). Unlike earlier clinical encoders trained on narrow or single-institution datasets, this model uses data from multiple institutions, domains, and regions. This broader training data improves generalization and enables the model to achieve state-of-the-art performance across many biomedical NLP benchmarks.
Beyond English biomedical text, ModernBERT has also been shown to be effective in non-English and highly technical clinical settings. In Japanese radiology report classification, ModernBERT achieved 39% reduction in training time and up to a 1.651.65× speed-up during training and inference, while maintaining comparable classification performance (Yamagishi2025ModernBERTRadiology).
These findings show that ModernBERT is well-suited for domain-specific applications that involve processing long documents efficiently, and specialized vocabularies. They act as motivation to use ModernBERT as a strong base model for further domain-adapted encoder development.
2.4.3 Architectural Analysis and Comparisons
In a controlled study designed to isolate architectural effects from differences in pre-training data, ModernBERT was compared with DeBERTaV3 when both models were pre-trained on the same dataset. The results indicate that while ModernBERT offers faster training and inference speeds attributable to its architectural optimizations, the earlier DeBERTaV3 architecture maintained superior sample efficiency and overall benchmark performance under identical pre-training conditions (Antoun2025ModernBERT). This finding underscores the importance of disentangling architectural improvements from data effects when evaluating advances in encoder design and suggests that modern architectural enhancements may need to be paired with training methodology innovations to fully realize performance gains.
Another line of architectural comparison came from NeoBERT (Breton2025NeoBERT), a next-generation bidirectional encoder that incorporates state-of-the-art architectural refinements, modern pre-training regimes, and optimized depth-to-width design. NeoBERT achieves state-of-the-art results on large evaluation suites such as GLUE and the Massive Text Embedding Benchmark (MTEB), regardless of its relatively small model size of 250250M parameters and extended context length. It outperforms standard baselines, including BERT Large, RoBERTa Large, NomicBERT, and ModernBERT under identical fine-tuning conditions (Breton2025NeoBERT). This comparison highlights how architectural choices such as optimal layer configurations, positional encoding mechanisms, and more balanced model scaling can yield efficiency and performance advantages even compared to models with greater nominal capacity.
Together, these comparative analyses illustrate active research interest in the architectural evolution of encoder-only transformers, where improvements are evaluated not only relative to legacy models like BERT and RoBERTa, but also against contemporary designs such as DeBERTaV3 and NeoBERT. Such work highlights trade-offs between efficiency, sample efficiency, and task performance. Moreover, it motivates ongoing investigation into how design choices impact downstream effectiveness across varied NLP benchmarks.
2.5 Turkish NLP Benchmarks
Early evaluation of Turkish NLP models has predominantly relied on individual, task-specific datasets, developed independently for problems such as named entity recognition (NER), sentiment analysis, dependency parsing, or question answering. Examples include Turkish NER datasets derived from Wikipedia (Rahimi2019WikiANN) and news (Tur2003MilliyetNER), sentiment datasets collected from the web (Amasyali2018Sentiment), and syntactic resources such as Universal Dependencies Turkish treebanks (UD25Turkish; UDTurkishIMST; UDTurkishBOUN). Even though these datasets have been useful in enabling a steady and consistent progress on task-specific problems, they often differ substantially in annotation quality, domain coverage, and evaluation protocols, and many lack widely adopted, fixed train/validation/test splits, limiting the comparability of reported results across studies.
As transformer-based models became dominant, evaluation practices for Turkish largely continued to follow this fragmented paradigm. Recent works frequently assembled their own evaluation datasets or introduced task-specific splits tailored to a single study. For instance, YTU-Cosmos-BERT evaluated models on tasks such as sentiment analysis, news classification, mask prediction, and zero-shot classification using custom-constructed datasets and internally defined splits, rather than established benchmark splits (Kesgin2023YTU). TurkishBERTweet focused on sentiment analysis and hate speech detection in social media, similarly relying on task-specific tweet datasets and cross-validation-based evaluation protocols rather than standardized benchmark test sets (Najafi2023TurkishBERTweet). While the stand-alone evaluations conducted within such works demonstrate practical effectiveness, they make it difficult to draw robust, reproducible comparisons between models trained under different conditions.
In response to these limitations, several Turkish NLP benchmark efforts have emerged that aim to consolidate multiple tasks under unified evaluation frameworks. Mukayese aggregated diverse Turkish NLP tasks, including language modeling, machine translation, NER, summarization, and text classification into a single benchmark suite, providing common baselines (Safaya2022Mukayese). TrGLUE adapted the GLUE paradigm to Turkish natural language understanding tasks, though its scope remained largely limited to sentence-level NLU (Altinok2025TrGLUE). More recently, Cetvel, a large-scale LLM-oriented evaluation benchmark, covered a broader set of discriminative and generative tasks while emphasizing cultural and linguistic relevance for Turkish (kuisai2024cetvel).
Despite recent advances, a systematic quality assessment of widely used Turkish benchmark datasets showed that a substantial portion of these datasets fail to meet basic criteria for linguistic correctness, coherence, and cultural appropriateness, underscoring the need for stricter dataset construction and validation standards (Umutlu2025Evaluating). Last but not least, most existing benchmarks focus on traditional NLP tasks, with limited coverage of modern NLP tasks such as academic text understanding, retrieval, and code-related tasks, which have become central to contemporary transformer and LLM evaluation.
Taken together, these observations motivate the need for a standardized Turkish NLP benchmark tailored to transformer-encoder models, characterized by (i) high-quality and well-documented datasets, (ii) fixed and community-accepted evaluation splits, (iii) diverse task coverage, including modern tasks such as coding, academic understanding, and retrieval. Such a benchmark would address the fragmentation evident in recent Turkish BERT studies (Kesgin2023YTU; Najafi2023TurkishBERTweet) and enable more reliable, reproducible, and comprehensive assessment of Turkish language models.
3 Data
3.1 Pre-training Data
We compiled a comprehensive dataset primarily consisting of Turkish monolingual text, supplemented with a small portion of English, coding and math data, to pre-train our model. To address significant size disparities and differences in quality across corpora and ensure balanced representation of different text types, we employed strategic oversampling of smaller but high-quality datasets. After applying the oversampling strategy, our final pre-training corpus consisted of 86.5886.58B tokens, of which 1313% were non-Turkish sources. In the rest of the paper, all references to the size of the pre-training corpus will include the effects of oversampling. The details of each corpus are explained in the following subsections, and the training corpora statistics, as well as the sampling rates, are summarized in Table 2.
Web Corpora
We extracted the tur_Latn subset of the FineWeb-2 (Penedo2025FineWeb2) as the main Turkish web corpus, with a knowledge cutoff date of April 2024. We also included a sample of 5 million English web pages from the FineWeb-2 English dataset to provide English language context during pre-training. Moreover, we included a Turkish Wikipedia dataset comprising articles scraped from Wikipedia’s Turkish-language pages in August 2024. This dataset provides encyclopedic knowledge covering a wide range of topics written in a formal and informative style.
Scientific Corpora
We used the Dergipark and Yöktez datasets from TURNA (Uludogan2024TURNA). Dergipark is a corpus of Turkish academic articles, and Yöktez is a corpus of Turkish graduate theses.
Book Corpus
We used the Book Corpus dataset from TURNA (Uludogan2024TURNA), a compilation of Turkish fiction and non-fiction books.
Bilkent Creative Writings
The Bilkent Creative Writings (Yilmaz2025BilkentWritings) corpus comprises short stories written by Bilkent University students while taking creative writing courses in Turkish.
ParlaMintTR
The ParlaMintTR corpus originates from the CLARIN Flagship project222clarin.eu/parlamint, and contains transcripts of Turkish parliamentary debates within the broader European ParlaMint initiative.
Code Corpus
The Code corpus consists of English code repositories, documentation and instructions obtained by randomly selecting a subset of 2.52.5M examples from several coding corpora (CodeBagel2024; GitHubCode2022; PythonCodeAlpaca2023; TinyCodes2023).
Math Corpus
The Math corpus comprises mathematical questions and solutions in English, collected from well-known academic and public datasets (Amini2019MathQA; HendrycksMath2021; Gao2024OmniMath; Yue2023Mammoth; Yu2023MetaMath; Hendrycks2024CompetitionMathNA).
Corpus Type # Docs # Tokens (B) Sampling Rate
FineWeb-2 Turkish Web $88,769,907$ $56.041$ 11
FineWeb-2 English Web $5,000,000$ $5.676$ 11
Wikipedia-TR Web $604,716$ $0.169$ 22
Dergipark Scientific $334,429$ $1.738$ 11
Yöktez Scientific $475,817$ $14.887$ 11
Books Literary $5,078$ $0.621$ 33
Bilkent Creative Writings Creative Text $8,457$ $0.007$ 33
ParlaMint-TR Dialogue $1,333$ $0.068$ 55
Code Code $2,500,000$ $5.415$ 11
Math Math $817,392$ $0.261$ 11
Table 2: Training Corpora Statistics: The complete pre-training corpus consists of 86.5886.58B tokens.
3.2 Fine-tuning Data
For fine-tuning, we introduce TabiBench, a unified evaluation benchmark meticulously crafted for Turkish NLP. TabiBench encompasses 28 datasets spanning 8 major task categories: text classification, token classification (covering named entity recognition and part-of-speech tagging), semantic textual similarity, natural language inference, question answering, information retrieval, code retrieval, and academic domain challenges. Our benchmark brings together carefully curated 21 Turkish datasets alongside 7 newly created machine-translated datasets, specifically for code retrieval and academic domain evaluation. Every dataset included in TabiBench undergoes a systematic quality assessment and is released with standardized train-validation-test splits to ensure reproducibility and enable fair comparisons across models. Further details on our evaluation datasets and the TabiBench framework are provided in Section 4.4.
4 Methodology
This section elaborates the methodology used in this project, including the tokenizer, model, pre-training objectives, fine-tuning and evaluation.
4.1 Tokenizer
In this work, we use the tokenizer developed for Kumru LLM (turker2025kumru).
Specifications
[…]
Training Dataset
The tokenizer was trained on a dataset mixture that consists of 95% Turkish and 5% English, code, and math—a similar mixture to our model’s pre-training data. This allows the models to perform well on both Turkish and English languages while supporting code and math.
Adaptation for TabiBERT
To use a tokenizer developed for a decoder-only LLM with TabiBERT, we added the special tokens [CLS] and [SEP] and modified the input processing template to make it compatible with BERT tasks.
Fertility
Fertility (Rust2021HowGood) is a common metric to measure the representation efficiency of a tokenizer, indicating the average number of tokens used to represent text. In a controlled setting where training data is assumed to be ample and diverse, vocabulary size is the main parameter that controls fertility. A large vocabulary allows lower fertility values at the cost of increased model size and, hence, computational and memory requirements (Ali2024Tokenizer). Although a lower fertility is desired in language modeling, there is a clear trade-off between vocabulary size and fertility.
In order to measure the representation efficiency of tokenizers used in this work, we measure fertility values by tokenizing 11M randomly selected web pages and report the relative fertility values compared to TabiBERT in Figure 1. It is observed that:
  • YTU-Cosmos-BERT provides the lowest fertility. This is the result of using an uncased tokenizer.
  • TabiBERT and BERTurk tokenizers, both being cased tokenizers, have comparable fertility values, despite TabiBERT tokenizer having a significantly larger vocabulary size. This is due to TabiBERT tokenizer’s pre-tokenization regex, which results in spending more tokens on numbers, punctuations and line-break characters. This is a conscious design choice and a natural trade-off to better represent these symbols.
  • TurkishBERTweet has a similar fertility value to TabiBERT and BERTurk, despite having a vocabulary size that is thrice and twice that of TabiBERT and BERTurk  respectively, and being an uncased tokenizer. We hypothesize that this is because it was trained and optimized for social media texts compared to other tokenizers that were trained on more general corpora.
  • Lastly, despite its large vocabulary of $256,000$ tokens, the mmBERT tokenizer’s fertility is significantly higher than that of Turkish tokenizers. This is due to Turkish being a relatively low-resource language and having a relatively small amount of data compared to other languages in corpora such as FineWeb-2 (Penedo2025FineWeb2).
[…]
Figure 1: Tokenizer fertility values of established Turkish BERT models and the modern multilingual mmBERT, relative to the proposed TabiBERT (represented by the dashed baseline). X axis labels indicate models along with the vocabulary sizes in parentheses. Y axis indicates relative number of tokens (fertility) required to represent Turkish texts compared to TabiBERT. Despite its large vocabulary of $256,000$ tokens, mmBERT suffers a 41.0% increase in fertility compared to TabiBERT in Turkish texts This demonstrates that despite architectural improvements, the lack of language-specific tokenizer optimization in mmBERT results in substantial subword fragmentation compared to TabiBERT.
4.2 Model
Our model is built on the ModernBERT (Warner2024ModernBERT) architecture. We use the ModernBERT-base setup, with 22 hidden layers, 12 attention heads, an intermediate size of 1,152, a hidden size of 768, and GLU (Shazeer2020Glu) expansion to 2,304. We modify the vocabulary size config to $50,176$ to match with the tokenizer’s vocabulary size.
Consequently, our model incorporates several recent improvements, including Rotary Positional Embeddings (RoPE) (Su2024Roformer) to support longer contexts, Local-Global Alternating Attention (Team2024Gemma) for efficient processing of long texts, as well as unpadding (Portes2023Mosaicbert; Zhang2024Mgte) and Flash Attention (Dao2022FlashAttention) mechanisms to enable faster inference and training.
Its inference speed is comparable to other models for a standard context length of 512; however, it is $2.65$ times faster than others when processing long contexts of $8,192$ tokens  (Warner2024ModernBERT). Moreover, for both short and long context lengths, it is can process inputs with twice the batch size thanks to its lower GPU memory footprint.
4.3 Pre-training Objectives
For pre-training, we adopt the Masked Language Modeling (MLM) objective as in ModernBERT, with a 30% token masking rate. We utilize the official ModernBERT pre-training codebase and recipe to ensure methodological consistency and reproducibility. The pre-training process is organized into 3 progressive phases, each exposing the model to 850B, 125B, and 25B tokens, respectively. Hence, we scale the ModernBERT pre-training recipe by 50% with the aim of pre-training for 11T tokens in total.
To best utilize the dataset and maximize training throughput, we pre-tokenize the entire pre-training corpus and chunk it into pieces that fit within the context length. We perform optimization using the StableAdamW algorithm(Wortsman2023StableLowPrecision), paired with a trapezoidal learning rate schedule (Xing2018Walk) and 1-sqrt decay (Hagele2024Scaling) for improved stability during large-scale training. The pre-training data is a mixture of samples from the datasets described in Section 3 and consists of 86.5886.58B tokens, of which 13% are non-Turkish. We pre-train our model for $117$ hours on a cluster of 8 x NVIDIA H100 GPUs, resulting in 11T tokens being exposed to the model, that is, $11.83$ epochs over the entire dataset.
4.4 TabiBench Evaluation Framework
4.4.1 Motivation and Design Principles
The evaluation of Turkish encoder models is constrained by several systemic shortcomings in existing benchmarks. It has long relied on small, task-specific datasets that are repeatedly reused across studies, increasing the risk of test set contamination and overfitting to benchmark idiosyncrasies (Kilinc2017; Yang2023Contamination). In addition, many commonly used resources lack explicit probing of Turkish-specific linguistic phenomena such as morphology and agglutination (Oflazer2018), and do not provide fixed, community-accepted train/validation/test splits, undermining reproducibility and comparability. Recent Turkish BERT studies further illustrate this fragmentation, frequently introducing ad-hoc evaluation datasets or custom splits tailored to individual works (Kesgin2023YTU; Najafi2023TurkishBERTweet). Although emerging benchmark efforts improve task consolidation, they remain limited either in task diversity or in evaluation rigor, and meta-analyses reveal persistent issues in dataset quality, annotation consistency, and cultural adequacy (Umutlu2025Evaluating).
These limitations motivate the development of TabiBench, a standardized evaluation framework explicitly tailored to Turkish transformer-encoder models. TabiBench is designed around three core principles: (i) high-quality, carefully validated datasets with clear documentation and linguistically informed construction, (ii) fixed and publicly released evaluation splits to support reproducible and fair model comparison, and (iii) broad task coverage that extends beyond classical classification and sequence labeling to include modern evaluation dimensions such as academic text understanding, retrieval-oriented reasoning, and code-related tasks. By addressing the fragmentation and quality concerns evident in prior Turkish NLP benchmarks, TabiBench aims to enable more reliable, comprehensive, and future-facing evaluation of Turkish encoder models.
  1. 1.
    Quality over quantity: All datasets undergo systematic review for annotation consistency, label reliability, and task validity
  2. 2.
    Standardization: Unified train-validation-test splits with explicit splitting methodology for reproducibility
  3. 3.
    Task diversity: Coverage of traditional NLU tasks alongside modern evaluation paradigms (retrieval, code understanding, domain specialization)
4.4.2 Dataset Selection and Quality Assessment
We systematically review existing Turkish fine-tuning datasets and analyze both dataset quality and task relevance for assessing model capabilities. Datasets are selected based on three mandatory criteria:
  1. 1.
    Clear pattern-label correspondence: The relationship between input text and output labels must be unambiguous and learnable
  2. 2.
    Mutually exclusive labels: Each example must belong to exactly one class, ensuring well-defined classification boundaries
  3. 3.
    Documented annotation guidelines: Clear, consistently applied annotation protocols must be available to guarantee label reliability
This manual quality assessment ensures our evaluation benchmarks reflect realistic and meaningful downstream applications rather than artifacts of poor annotation or ambiguous task definitions.
4.4.3 Standardized Split Generation Methodology
A significant challenge in Turkish NLP evaluation is the absence of standardized dataset splits. Many existing datasets lack explicit train-validation-test divisions, leading researchers to create inconsistent splits that prevent fair cross-study comparisons. We address this through a systematic split generation methodology applied uniformly across all datasets:
  • If a train-validation-test split is available, we use the existing divisions as provided.
  • For datasets with a train-test split only, we create a validation split from the training set, sized to match the test set, and apply this across all models.
  • In cases with a train-validation split, we reassign the validation set as the test split, then generate a new validation split from the training data following the approach above.
  • In cases with a validation-test split, we split validation into train and validation sets in 80% and 20% proportions, respectively.
  • When only a single combined split is present, we partition the data into train, validation, and test sets in 70%, 15%, and 15% proportions, respectively.
All splits use stratified sampling to preserve class distributions and ensure statistical validity. To support reproducibility and community adoption, we release these standardized evaluation suites publicly on HuggingFace (TabiBench).
TabiBench encompasses diverse task categories, combining curated existing Turkish datasets with newly created resources.
4.4.4 Task Categories and Dataset Coverage
Our fine-tuning and evaluation benchmark covers the following task categories and corresponding Turkish datasets:
  • Text classification includes: NewsCat (McemilgNewsCat), which contains news articles categorized into five classes; Product Reviews (FthbrmnbyTurkishProductReviews), a binary sentiment classification dataset; Bil Tweet News (Toraman2017EventPrediction; Toraman2021EventRetrieval), a four-class news classification dataset based on social media posts; and Gender Hate Speech Turkish (Toraman2022HateSpeech; Sahinuc2023GenderBias), a three-class dataset for detecting gender-based hate speech.
  • Token classification includes: WikiNER (Altinok2023Turkish) and WikiANN-TR (Rahimi2019WikiANN), which provide Turkish named entity recognition annotations; and POS-UD BOUN (UDTurkishBOUN) and POS-UD IMST (UDTurkishIMST), which provide Turkish part-of-speech annotations based on Universal Dependencies.
  • Semantic textual similarity tasks include: SICK-TR (Dehghan2025Turkish) and STSb-TR (BekenFikri2021Summarization), which provide Turkish sentence pairs annotated with semantic similarity scores.
  • Natural language inference tasks include: SNLI-TR and MultiNLI-TR (BudurEtAl2020Data), which are derived from the Natural Language Inference in Turkish (NLI-TR) dataset.
  • Question answering tasks include: TQuAD (ErdometoTQuad2) and XQuAD (Artetxe2019CrossLingual), which provide Turkish datasets for reading comprehension and cross-lingual question answering.
  • Information retrieval tasks follow the Turkish translations provided by the TR-MTEB initiative (BaysanGungor2025TR-MTEB) and include: BiText, MsMarco-TR, Scifact-TR, Fiqa-TR, NFCorpus-TR, and Quora-TR. These datasets cover diverse domains such as general web search, scientific literature, financial documents, and question–answer retrieval.
  • Code retrieval tasks are translated from English using GPT-4.1 while preserving code integrity:
    • Apps-Retrieval (Hendrycks2021Measuring) consists of programming problem descriptions as queries and corresponding code solutions as documents. Its Turkish variant, Apps-TR, is created by translating only the problem descriptions.
    • CodeSearchNet (Husain2019CodeSearchNet) contains code snippets paired with docstrings across six programming languages. The Turkish variant, CodeSearchNet-21K-TR, is constructed by selecting $2,500$, $500$, and $500$ examples per programming language for training, validation, and testing, respectively, and translating the docstrings.
    • CosQA (Huang2021CosQA) contains programming questions paired with corresponding code solutions. The Turkish version, CosQA-TR, is created by translating the questions while keeping the code unchanged.
    • StackOverflow-QA (Li2025Coir) consists of programming questions paired with their highest-voted answers. The Turkish version, StackOverflowQA-TR, is obtained by translating the natural language components of both questions and answers.
  • Academic domain tasks assess model performance on specialized scholarly text. Three datasets are translated from English using GPT-4.1, while one dataset is natively constructed in Turkish:
    • MedNLI (Romanov2018Lessons) is a medical-domain natural language inference dataset consisting of sentence pairs annotated with entailment labels.
    • PubMedRCT-20K (Dernoncourt2017Pubmed) is a sentence-level classification dataset where sentences from scientific abstracts are labeled with rhetorical roles. A balanced subset of $10,000$ examples is selected using stratified sampling before translation.
    • SciCite (Cohan2019Structural) is a three-class citation intent classification dataset, labeling citations as background, method, or result.
    • Thesis-Abstract-Classification-11K is a natively constructed Turkish dataset derived from thesis abstracts collected from the National Thesis Center of Turkey (TurkishThesis2025Ertugrul). The task is to predict the field of study of a thesis abstract among 187 categories. A balanced subset with 60 samples per category is constructed, making this a challenging multi-class classification task due to the large label space and specialized academic language.
5 Evaluation
This section presents a comprehensive evaluation of TabiBERT across a diverse range of Turkish NLP tasks from TabiBench. We first describe the models compared in our evaluation, including state-of-the-art monolingual Turkish baselines and the multilingual mmBERT model. We then detail our evaluation methodology, including task and dataset selection, hyperparameter tuning procedures, fine-tuning configurations, context length selection, training objectives, and evaluation metrics. Finally, we present detailed results across eight task categories, demonstrating TabiBERT’s performance relative to baseline models and analyzing key strengths and limitations.
We begin by summarizing the key results:
  • TabiBERT achieves state-of-the-art performance among Turkish models, outperforming all monolingual baselines with a total average score of $77.58$, surpassing the previous best model by $1.62$ points.
  • The largest improvement is observed in question answering tasks, where TabiBERT achieves an F1 score of $69.71$, outperforming the next best Turkish model by $9.55$ points (1616% relative improvement).
  • TabiBERT leads the Turkish models in five out of eight task categories, including code retrieval where it outperforms the previous best by $2.41$ points, and information retrieval where it achieves an NDCG@10 score of $75.44$.
5.1 Models Compared
We compare TabiBERT with three Turkish BERT models, configurations of which are based on BERT-Base, representing the monolingual state-of-the-art approaches for Turkish. They are BERTurk, YTU-Cosmos-BERT, and TurkishBERTweet. YTU-Cosmos-BERT and TurkishBERTweet are selected on the fact that they have a clear recipe in the form of publications. Even though BERTurk is not the product of an academic work, it is a well-established and widely used model in Turkish NLP research as a strong baseline. Moreover, all of the mentioned models have a large, diverse pre-training corpus. In addition to Turkish BERT models, we also compare against the recently introduced ModernBERT-based multilingual encoder mmBERT.
BERTurk
(Schweter2020BERTurk) is a widely used Turkish monolingual encoder-only model, pre-trained on the Web and special corpora. It serves as a strong baseline for Turkish NLP tasks and has influenced much of the subsequent Turkish NLP research.
YTU-Cosmos-BERT
(Kesgin2023YTU) is an uncased Turkish BERT model, pre-trained on a large Turkish corpus including The Web, Wikipedia and books.
TurkishBERTweet
(Najafi2023TurkishBERTweet) TurkishBERTweet is another uncased Turkish BERT model, pre-trained in large scale, tailored for social media.
mmBERT
(Marone2025mmBERT) mmBERT is a multilingual encoder based on ModernBERT, with the aim of obtaining a multilingual ModernBERT model.
5.2 Methodology
We conduct a comprehensive evaluation of 5 models across various Turkish NLP tasks, combined under the TabiBench benchmark. To ensure fair and reproducible comparisons, we establish an end-to-end pipeline for task selection, dataset preparation, hyperparameter tuning, and fine-tuning.
Task and Dataset Selection
Our evaluation includes a diverse set of tasks that assess model capabilities across various domains and applications. These tasks include natural language understanding tasks such as text classification, natural language inference, semantic textual similarity, named entity recognition, and part-of-speech tagging, as well as more specialized tasks, including question answering, information retrieval, code retrieval, and academic domain tasks. We develop TabiBench, a comprehensive benchmark consisting of twenty eight datasets spanning these eight task types, as further elaborated in Section 4.4.
Hyperparameter Tuning
We perform systematic hyperparameter tuning for all model-task pairs to ensure optimal performance. We determine our hyperparameter search space based on prior work (He2023DeBERTaV3; Warner2024ModernBERT; Breton2025NeoBERT) and our compute budget, ensuring effective fine-tuning across all tasks and models. Our hyperparameter search space is presented in Table 3.
Table 3: Hyperparameter Search Space
Parameter Values
Learning Rate $5\text{\times}{10}^{-6}$, $1\text{\times}{10}^{-5}$, $2\text{\times}{10}^{-5}$, $3\text{\times}{10}^{-5}$
Weight Decay $\text{\times}{10}^{-5}, 1×1$$\text{\times}{10}^{-6}$
Batch Size 16, 32
Epoch 10, with early stopping
Early Stopping
All models are trained for up to 10 epochs with early stopping, with a patience of 2 epochs, based on validation split loss to prevent overfitting. The best hyperparameter configuration for each model-task pair is determined through validation set performance, and these configurations are then used for final fine-tuning and test set evaluation. For the uncased YTU-Cosmos-BERT and TurkishBERTweet models, lower-case conversion is applied to avoid problems with the Turkish character "I":
text = text.replace("I", "ı").lower()
Fine-tuning Procedure
For each model-task pair, we fine-tune using the optimal hyperparameters identified during tuning. All fine-tuning experiments use a linear decay with warmup scheduler, where the learning rate warms up to the full value over 6% of the training duration and linearly decays to 0.02x the full learning rate by the end of training. We employ mixed precision training and set a random seed of 25 for reproducibility across all experiments. These configurations are based on the example finetuning configs released in the ModernBERT repository 333github.com/AnswerDotAI/ModernBERT/yamls/finetuning/hf-bert-base-uncased.yaml. For each fine-tuning task category, we report a single score by computing a weighted average across all datasets in that category, where each dataset’s weight is proportional to its test set size. This methodology ensures that larger, more representative datasets have a corresponding influence on the overall results, due to the variation in test set sizes ranging from 150 to $35,000$ examples.
Context Length
To determine appropriate context lengths for fine-tuning, we compute the 95th percentile of sequence lengths for each dataset. For each model-task pair, we set the maximum context length to the minimum of the model’s context length and the dataset’s 95th percentile length, making sure that no sequences are truncated unnecessarily while respecting each model’s architectural constraints. Since all other Turkish models have a context length of $512$ tokens, whereas TabiBERT and mmBERT support up to $8,192$ tokens, the latter two are able to utilize longer context windows for tasks where the 95th percentile exceeds $512$ tokens, providing them with an advantage on longer sequences.
Training Objectives
We apply task-specific training objectives for fine-tuning across different task types. For text classification and natural language inference tasks, we use cross-entropy loss, implemented through HuggingFace’s default loss computation for sequence classification models. Token classification tasks, including named entity recognition and part-of-speech tagging, are trained using cross-entropy loss applied at the token level. Semantic textual similarity tasks are formulated as regression problems, using a custom regression model wrapper that optimizes mean squared error (MSE) loss to predict continuous similarity scores. Question answering tasks employ span prediction, where the model predicts the start and end positions of answer spans within the context using cross-entropy loss for each position. For information retrieval and code retrieval tasks, we use contrastive learning objectives to learn representations that maximize the similarity between queries and relevant documents while minimizing similarity with irrelevant documents.
Evaluation Metrics
We use task-appropriate evaluation metrics: macro-averaged F1 for classification and NLI (to account for class imbalance), micro-averaged F1 at the word level for token classification, Pearson correlation for semantic similarity, F1 for question answering, and NDCG@10 for information retrieval and code retrieval tasks.
5.3 Results
The performance of TabiBERT compared to monolingual Turkish models and mmBERT across all tasks is presented in Table 4. TabiBERT outperforms all other Turkish models in total average score across all tasks, achieving a score of 77.58, surpassing the previous best Turkish model, BERTurk, by 1.62 points. Moreover, the improvement is consistent across tasks with TabiBERT achieving the highest score in 5 out of 8 task categories, and being competitive in the remaining tasks. Despite being optimized for long context, TabiBERT maintains strong performance across tasks, demonstrating its versatility and effectiveness across a wide range of applications.
Table 4: Comparison of downstream task performance across all evaluated models. For each column, the highest score among the Turkish models (excluding the multilingual mmBERT) is shown in bold. The evaluation metric used for each task type is also displayed in the column headers. TabiBERT achieves the highest score in five out of eight task categories among Turkish models.
Model # of params Text Clf Token Clf STS NLI QA Academic Retrieval Code Retrieval Total Avg
M F1 F1 Pearson F1 F1 F1 NDCG@10 NDCG@10 (TabiBench)
TurkishBERTweet $163$ $79.71$ $92.02$ $75.86$ $79.10$ $38.13$ $63.12$ $68.40$ $43.49$ $67.48$
YTU-Cosmos-BERT $111$ 84.25 $93.60$ $84.68$ $84.16$ $31.50$ $71.78$ $74.29$ $53.80$ $72.26$
BERTurk $110$ $83.42$ 93.67 85.33 $84.33$ $60.16$ $71.40$ $74.84$ $54.54$ $75.96$
TabiBERT $149$ $83.44$ $93.42$ $84.74$ 84.51 69.71 72.44 75.44 56.95 77.58
mmBERT $307$ $82.54$ $93.81$ $87.05$ $84.38$ $71.47$ $72.65$ $76.20$ $66.02$ $79.26$
Core Natural Language Understanding Tasks
Being always either on par with or slightly above the strongest Turkish baselines, TabiBERT demonstrates strong performance across core natural language understanding tasks: text classification, token classification, semantic textual similarity, and natural language inference. Text classification results of all models except for TurkishBERTweet are relatively close to each other, with YTU-Cosmos-BERT achieving the highest score of 84.25, followed closely by TabiBERT with 83.44. All Turkish models achieve high scores above 92 in token classification tasks with BERTurk leading with 93.67, closely followed by YTU-Cosmos-BERT and TabiBERT with 93.60 and 93.42, respectively. As it is in the case of text classification, all models except for TurkishBERTweet achieve similar scores in semantic textual similarity tasks with BERTurk leading with 85.33, followed by TabiBERT with 84.74. TabiBERT takes the lead in natural language inference tasks with a score of 84.51, slightly surpassing BERTurk’s score of 84.33 and YTU-Cosmos-BERT’s score of 84.16. In STS and NLI tasks, TabiBERT consistently matches or marginally exceeds BERTurk and YTU-Cosmos-BERT, demonstrating that the native long context of TabiBERT does not come at the expense of sentence-level semantic precision. The close scores across the top three models indicate that these tasks are close to saturation.
Question Answering
The largest margin of improvement for TabiBERT over other Turkish models is observed in question answering tasks. TabiBERT outperforms all monolingual baselines by a considerable margin, achieving an F1 score of 69.71 compared to the next best score of 60.16 by BERTurk. There are two factors in explaining this large margin: 1) TabiBERT excels at more difficult tasks. 2) TabiBERT’s long context length allows it to process long contexts in the TQuAD dataset without information loss to truncation.
Academic Domain Tasks
Performance gains are also evident in specialized academic tasks, including NLI and multi-class text classification. Trained on a corpus enriched with Turkish scientific articles and graduate theses, TabiBERT achieves the highest score among Turkish models, with particularly strong results on Thesis-Abstract-Classification-11K, a challenging long context multi-class task with limited per-class data. Thus, TabiBERT’s pre-training on academic texts and its long context capabilities contribute to its superior performance in this domain.
Retrieval and Code Retrieval
TabiBERT outperforms all other Turkish models in information retrieval and code retrieval tasks. TabiBERT consistently surpasses prior monolingual models on TRMTEB retrieval benchmarks, achieving an NDCG@10 score of 75.44 and surpassing the 74.84 score of the previous best Turkish model, BERTurk. The improvement is even more evident in code retrieval tasks, where TabiBERT achieves an NDCG@10 score of 56.95, considerably outperforming BERTurk by 2.41 points and YTU-Cosmos-BERT by 3.15 points. The benefits of extended context length and exposure to heterogeneous pre-training data, including bilingual text and code snippets, are particularly pronounced in retrieval and code retrieval tasks.
Comparison with Multilingual mmBERT
While TabiBERT outperforms all monolingual Turkish models, it still lags behind the multilingual mmBERT model, which achieves a total average score of 79.26. The explanation is threefold:
  • mmBERT is a larger model with 307307M parameters, compared to TabiBERT with 149149M parameters.
  • mmBERT’s tokenizer spends 41% more tokens than TabiBERT’s tokenizer to represent Turkish texts.
  • mmBERT is pre-trained on a massive multilingual corpus for a longer duration.
Although mmBERT still follows the BERT-base configuration with the same number of non-embedding parameters, despite its higher memory requirements, its large embedding matrix provides it with greater capacity to learn complex patterns. Moreover, mmBERT’s 41% higher fertility 4.1 results in greater fragmentation of Turkish texts, allowing it to better capture the nuances of Turkish morphology, as the tokenization process breaks words into smaller character ngrams. Although this comes at a cost, it allows mmBERT to perform better by allocating more compute on a given text. mmBERT’s pre-training on a massive multilingual corpus of $1,800$ languages, including high-resource languages such as English, enables it to learn richer representations by leveraging knowledge across languages. Not only that, but it is also pre-trained for a total of 33T tokens, significantly more than TabiBERT’s 11T tokens, which has been only possible by conducting more than $11.83$ epochs over the relatively small Turkish corpus.
Despite its strong performance across Turkish NLP tasks, mmBERT’s large size and multilingual nature may introduce inefficiencies for applications focused solely on Turkish. As a result of its large embedding matrix and the total number of parameters, mmBERT has significantly higher memory requirements than TabiBERT. Moreover, although both models have a native context length of $8,192$ tokens, mmBERT tokenizer’s high fertility results in increased token counts for Turkish texts and a 29% reduction in effective context length compared to TabiBERT.
Comparison Against Task-Specific Best Models
TabiBERT remains competitive when compared against the best-performing model for each individual task, as shown in Table 5, It achieves an average score of $77.58$, surpassing the task-specific best models’ average of $76.11$ by $1.47$ points. Although this improvement is slightlysmaller than in the comparison against all models, TabiBERT still outperforms the task-specific best in 5 of the 8 task categories, with particularly strong improvements in question answering ($9.55$ points), code retrieval ($2.41$ points), and academic domain tasks ($0.66$ points). For the remaining tasks where TabiBERT does not exceed the task-specific best, the performance differences are minimal, ranging from 0.81-0.81 to 0.25-0.25 points, indicating that TabiBERT remains highly competitive even against the strongest models for individual tasks.
Table 5: TabiBERT performance compared to the previous best-performing Turkish model on each downstream task. For each task, we report TabiBERT’s score, the previous best model’s score and name, and the absolute difference ($\Delta$). Boldfaced values indicate the best score per task.
Task TabiBERT Prev. Best Model $\Delta$
Text Clf $83.44$ $84.25$ YTU-Cosmos-BERT 0.81-0.81
Token Clf $93.42$ $93.67$ BERTurk 0.25-0.25
STS $84.74$ $85.33$ BERTurk 0.59-0.59
NLI $84.51$ $84.33$ BERTurk $0.18$
QA $69.71$ $60.16$ BERTurk $9.55$
Academic $72.44$ $71.78$ YTU-Cosmos-BERT $0.66$
Retrieval $75.44$ $74.84$ BERTurk $0.60$
Code Retrieval $56.95$ $54.54$ BERTurk $2.41$
Average $77.58$ $76.11$ $1.47$
5.3.1 Challenges and Adaptations
During our development and evaluation of TabiBERT, we encountered several challenges that required careful consideration and adaptation of our methodology.
Pre-training Throughput and Pre-tokenization
During our initial pre-training trials, we encountered extremely low throughput on 8×H100 SXM GPUs, achieving around only 3030% of the speed reported in the ModernBERT paper. We also ran training speed experiments on 8xA100SXM and training speed decreased only very slightly, which led us to believe that the bottleneck is the data loader and H100 GPUs are not utilized at their full potential. After consulting with the ModernBERT authors 444github.com/AnswerDotAI/ModernBERT/issues/236, we learned that the original authors applied pre-tokenization process to the training data to improve the throughput. This pre-tokenization step resolved the throughput issue, allowing us to achieve the training speeds reported in the ModernBERT paper. This experience highlights the critical role of thoroughly documented pre-training procedures in ensuring both reproducibility and scalability. With this lesson in mind, we mindfully documented the pre-training and fine-tuning processes for our own model to facilitate future research and practical adoption.
Evaluation Dataset Selection
During our evaluation, we initially selected the most commonly used supervised fine-tuning datasets in Turkish NLP research. However, after closer examination, we discovered that some of these datasets exhibited quality issues, including inconsistent annotations, ambiguous labeling criteria, or insufficient dataset size. To ensure reliable evaluation, we carried out a manual review of all candidate datasets, evaluating their annotation quality, label consistency, and task relevance. We selectively included only the higher-quality datasets in our evaluation suite, prioritizing those with clear annotation guidelines and reliable labels that are widely recognized in Turkish NLP research. This experience made evident the lack of standardized evaluation benchmarks in Turkish NLP research. To address this gap, we developed TabiBench, a comprehensive benchmark specifically designed for Turkish NLP, as further elaborated in Section 4.4.
Evaluation Methodology
For token classification tasks, including named entity recognition and part-of-speech tagging, we initially employed subword-level evaluation, which proved problematic. Different models tokenize the same word differently, with some tokenizers producing more subword tokens per word than others. This tokenization discrepancy made subword-level scores incomparable across models, as model tokenizers with higher token fertility would be able to be score higher by trivially predicting more fragmented subword tokens per word. To address this, we modified our evaluation methodology to use word-level evaluation instead, ensuring that predictions are properly aggregated at word boundaries before computing metrics. This approach provides a fair comparison across models regardless of their tokenization strategies.
These challenges highlight the importance of careful dataset selection, appropriate evaluation methodology design, and metric selection tailored to each task’s specific requirements.
6 Conclusion and Future Work
This work introduced TabiBERT and TabiBench. TabiBERT is a Turkish encoder model that leverages the state-of-the-art architectural innovations of the ModernBERT. It is pre-trained from scratch on a large, diverse corpus of native Turkish texts from the web, scientific, literary, academic, and parliamentary domains, complemented with limited bilingual, code, and math data. It achieves state-of-the-art results across a wide range of Turkish NLP tasks while being efficient in inference, with high speed and a small memory footprint. The gains from its 16 times longer context support compared to earlier models are demonstrated across multiple tasks without sacrificing performance on short-context tasks. The adoption of a modern tokenizer with explicit support for code, math, and structured text broadens the applicability of Turkish encoder models compared to prior ones.
TabiBench is a new comprehensive benchmark for Turkish NLP that encompasses a diverse set of tasks, including text classification, sequence labeling, question answering, and retrieval. It is designed to serve as a standard evaluation suite for Turkish language understanding models, with carefully curated high-quality datasets. Moreover, it covers a wide range of domains, including social media, product comments, treebanks, news, web search, academic and medical texts, and code. Every task in TabiBench includes standard train, validation, and test splits, along with clear evaluation metrics for consistent assessment.
Experiments conducted on TabiBench demonstrate the effectiveness of TabiBERT. It outperforms existing monolingual Turkish models across a broad range of tasks and establishes a new state of the art in overall benchmark performance. As a result of large-scale pre-training on huge, multilingual corpora and cross-lingual transfer, multilingual models such as mmBERT may achieve higher absolute scores. However, our results point out that a monolingual encoder whose design decisions were carefully made provides a strong trade-off between efficiency and performance on Turkish-centric applications. The results show that by utilizing modern encoder architectures, extended context lengths, and large-scale pre-training on large, high-quality, diverse and language-specific datasets, substantial gains are achieved for Turkish, known for being a morphologically rich and underrepresented language. Beyond absolute scores, this work provides a standardized, reproducible evaluation methodology for Turkish NLP, addressing long-standing issues of inconsistent dataset splits and fragmented evaluation practices.
We hope that TabiBERT and TabiBench will serve as valuable resources for the Turkish NLP community, fostering further research and development in this area.
There are several promising directions for extending this work. TabiBench already covers a wide range of tasks; however, it can be expanded to include more, such as math and long context tasks, enabling further evaluation of TabiBERT under these conditions. Furthermore, its long context support can be exploited in processing longer texts in tasks such as document classification and long-context question answering. Moreover, a general, state-of-the-art text embedding model can be mid-trained starting from TabiBERT to enhance performance and support longer contexts for retrieval tasks. Finally, pre-training a larger version of TabiBERT with a BERT-large configuration could be explored to potentially achieve even better results across tasks.
7 Limitations
Pre-training
Given our computational limitations, we were able to pre-train only the base model, comprising 149149M parameters, whereas the ModernBERT family also includes a larger 395395M parameter variant. In addition, our pre-training corpus of 86.5886.58B tokens is modest compared to large corpora such as FineWeb, consisting of 1818T tokens. Moreover, TabiBERT is pre-trained for 11T tokens, compared to ModernBERT and mmBERT which were pre-trained for 22T and 33T tokens respectively. As a result of limited Turkish corpora, TabiBERT was trained for $11.83$ epochs on the same data. It can be argued that this might limit its overall generalization and expose it to risks of overfitting or memorization. However, we hypothesize that the limited model size and consequently, capacity should mitigate these risks to great extent. Larger models and more diverse, expensive corpora could further enhance the linguistic capabilities and robustness of Turkish language models.
Fine-Tuning Experiments
Due to time and compute budget constraints, we were not able to conduct neither statistical significance testing, nor conduct fine-tuning experiments with multiple runs for each task using different random seeds, which prevents us from reporting more stable averages for downstream performance. Moreover, a portion of our benchmark consists of machine-translated datasets from English, specifically, all retrieval and code retrieval tasks, as well as three of four academic domain tasks. These translations, despite being carried out by a state-of-the-art LLM such as GPT-4.1, while useful for benchmarking and coverage, may not fully capture the intricacies of native Turkish.
Long Context Capabilities
While our model is capable of processing long input contexts, up to $8.192$ tokens, which is $16$ times of other major Turkish models, due to the lack of Turkish finetuning tasks with long-context, we cannot demonstrate the full advantage of this architectural feature at this time. Expanding available Turkish datasets to better reflect real-world, high-context use cases remains an important area for future work.
Data
Due to licensing restrictions on some source materials (academic publications, books), we cannot release the complete pre-training corpus. However, we provide detailed statistics, composition breakdowns, and preprocessing methodology to enable replication with alternative corpora.
8 Ethics Statement
Web content carries the risk of harmful content including bias, toxicity, abuse, and obscenity. In removing such harmful content from the training data, we rely on the methods used by the providers of the web corpus. Despite the filtering and cleaning efforts, there is still a risk that some harmful content remains in the training data. Although TabiBERT is primarily intended for use in text understanding tasks, the base model may still be used to generate text with the masked language modeling objective. Therefore, it is possible for TabiBERT to be affected by such harmful content and produce biased or toxic outputs. This calls for continuous monitoring of this system to eliminate such occurrences.
Heterogeneous sources such as academic texts, books, and parliamentary records were included in the training data to ensure linguistic coverage and representativeness for Turkish. Nevertheless, users should be mindful of ethical considerations related to data provenance and domain-specific bias when applying the model.
TabiBERT and TabiBench are released as research artifacts to support the development and evaluation of Turkish NLP systems. We encourage responsible use, task-specific evaluation, and bias-aware analysis when deploying the model in downstream applications.
Acknowledgments
We thank VNGRS [^5] https://vngrs.com/ for funding the computational resources used for model pre-training and fine-tuning. We also thank Lambda Labs [^6] https://lambda.ai/ for providing additional GPU resources. Finally, we thank Selman Baysan for his contributions to the implementation of the retrieval tasks.
8.1 AI Assistance
We acknowledge the use of AI assistants for grammar checking and language refinement during the preparation of this paper.
Appendix A Fine-Tuning Results
A.1 Fine-tuning Results
The appendix presents detailed performance comparisons across all evaluation tasks. Each table compares TabiBERT against established monolingual Turkish models with the best-performing monolingual model highlighted in each task. We also include mmBERT as a reference point. While the very large-scale multilingual model mmBERT often achieves competitive or superior performance, direct comparison is complicated by substantial differences in model scale, training data volume, and cross-lingual training objectives.
A.1.1 Text Classification
Model performance is evaluated across four text classification datasets: NewsCat (McemilgNewsCat), BilTweetNews (Toraman2017EventPrediction; Toraman2021EventRetrieval), GenderHateSpeech (Toraman2022HateSpeech; Sahinuc2023GenderBias), and ProductReviews (FthbrmnbyTurkishProductReviews). Table 6 presents these results using the macro F1 score. With the exception of TurkishBERTweet, the models achieve comparable results across the selected tasks. The YTU-Cosmos-BERT model yields the highest performance with a macro F1 score of $84.25$, followed by TabiBERT at $83.44$. Although TabiBERT does not attain the highest score on individual benchmarks, its is comparable with the top-performing models. Notably, mmBERT underperforms relative to the Turkish models with an average macro F1 score of $82.54$.
Table 6: Downstream performance of models on Text classification (eval metric: macro F1). The number of test samples per task is also given. For each column, the highest score among the Turkish models (excluding mmBERT) is shown in bold.
Model NewsCat BilTweetNews GenderHateSpeech ProductReviews Weighted Avg
Test Size $250$ $150$ $2,000$ $35,275$
TurkishBERTweet $91.98$ $53.16$ $68.58$ $80.37$ $79.71$
YTU-Cosmos-BERT $97.21$ $53.39$ $71.02$ $85.04$ $84.25$
BERTurk $95.60$ $57.87$ $68.25$ $84.30$ $83.42$
TabiBERT $95.20$ $50.11$ $69.01$ $84.32$ $83.44$
mmBERT $94.80$ $49.06$ $66.45$ $83.51$ $82.54$
A.1.2 Token Classification
Token classification performance is evaluated using four datasets: WikiNER (Altinok2023Turkish), WikiANN TR (Rahimi2019WikiANN), PosUDBOUN (UDTurkishBOUN), PosUDIMST (UDTurkishIMST). Table 7 presents the micro F1 scores. All Turkish models demonstrate high performance, with micro F1 scores exceeding $92$. The BERTurk model achieves the highest overall score of $93.67$ taking the leading position across all datasets. Consistent with the results observed in text classification tasks, TabiBERT maintains competitive performance levels.
Table 7: Downstream performance of models on Token classification (eval metric: micro F1). The number of test samples per task is also given. For each column, the highest score among the Turkish models (excluding mmBERT) is shown in bold.
Model WikiNER WikiANN TR PosUDBOUN PosUDIMST Weighted Avg
Test Size $1,000$ $10,000$ $979$ $1,100$
TurkishBERTweet $72.6$ $94.05$ $89.74$ $93.21$ $92.02$
YTU-Cosmos-BERT $78.96$ $95.41$ $89.19$ $94.30$ $93.60$
BERTurk $79.44$ $95.37$ $89.82$ $94.60$ $93.67$
TabiBERT $76.41$ $95.34$ $90.40$ $94.07$ $93.42$
mmBERT $75.97$ $95.85$ $90.72$ $94.26$ $93.81$
A.1.3 Semantic Textual Similarity
The performance in semantic textual similarity is evaluated with two datasets: SICK-TR (Dehghan2025Turkish) and STSb-TR (BekenFikri2021Summarization). Table 8 reports the results using the Pearson correlation score. All models, with the exception of TurkishBERTweet, achieve similar results with BERTurk performing the best with $85.33$. TabiBERT follows closely at $84.74$.
Table 8: Downstream performance of models on Semantic Textual Similarity (eval metric: Pearson correlation). The number of test samples per task is also given. For each column, the highest score among the Turkish models (excluding mmBERT) is shown in bold.
Model SICK-TR STSb-TR Weighted Avg
Test Size $4,927$ $1,379$
TurkishBERTweet $78.35$ $66.96$ $75.86$
YTU-Cosmos-BERT $85.27$ $82.55$ $84.68$
BERTurk $85.95$ $83.12$ $85.33$
TabiBERT $85.00$ $83.84$ $84.74$
mmBERT $87.81$ $84.31$ $87.05$
A.1.4 Natural Language Inference
Natural language inference capabilities are evaluated using two Turkish datasets: SNLI-TR and MultiNLI-TR (BudurEtAl2020Data). Table 9 reports model performance based on the macro F1 score. With the exception of TurkishBERTweet, the models exhibit similar performance levels with macro F1 scores near $84$. These results suggest that the leading models, including the multilingual mmBERT, are comparable for Turkish NLI tasks. Further performance improvements may necessitate more complex datasets or architectural innovations.
Table 9: Downstream performance of models on Natural Language Inference (eval metric: macro F1). The number of test samples per task is also given. For each column, the highest score among the Turkish models (excluding mmBERT) is shown in bold.
Model SNLI-TR MultiNLI-TR Weighted Avg
Test Size $9,824$ $4,923$
TurkishBERTweet $83.05$ $71.21$ $79.10$
YTU-Cosmos-BERT $87.18$ $78.15$ $84.16$
BERTurk $87.21$ $78.57$ $84.33$
TabiBERT $86.47$ $80.60$ $84.51$
mmBERT $86.44$ $80.28$ $84.38$
A.1.5 Question Answering
Question answering performance is evaluated using two Turkish datasets: TQuAD (ErdometoTQuad2) and XQuAD (Artetxe2019CrossLingual). Table 10 presents the F1 scores obtained by each model. Among all evaluated tasks, this benchmark reveals the most substantial performance gain achieved by TabiBERT over other Turkish models with an F1 score of $69.71$ in comparison to the next best result of $60.16$ recorded by BERTurk.
Two primary factors likely contribute to this marked improvement. First, the architectural enhancements of TabiBERT, specifically its deeper encoder and refined attention mechanism, appear effective for the span detection and contextual reasoning needed for question answering. Second, all models were finetuned using context lengths based on the 95th percentile of each dataset, specifically $1,024$ tokens for TQuAD and $833$ for XQuAD. Since the 95th percentile of TQuAD reaches $1,024$ tokens exceeding the $512$ maximum supported by the other Turkish models, only TabiBERT with its $8,192$ token capacity could be finetuned without truncation. Consequently, only TabiBERT could process the TQuAD examples without information loss during both training and inference.
As for XQuAD, limited volume of training data ($833$ training and only $179$ test examples) likely constrains model learning, resulting in lower scores across all models irrespective their context handling capabilities.
Table 10: Downstream performance of models on Question Answering (eval metric: F1). The number of test samples per task is also given. For each column, the highest score among the Turkish models (excluding mmBERT) is shown in bold.
Model Tquad XQuAD Weighted Avg
Test Size $2,520$ $179$
TurkishBERTweet $38.04$ $39.40$ $38.13$
YTU-Cosmos-BERT $32.01$ $24.25$ $31.50$
BERTurk $63.30$ $15.96$ $60.16$
TabiBERT $72.34$ $32.61$ $69.71$
mmBERT $71.55$ $70.40$ $71.47$
A.1.6 Information Retrieval
We evaluate information retrieval performance on six datasets: BiText, MsMarco-TR, Scifact-TR, Fiqa-TR, NFCorpus-TR, and Quora-TR, which are translated into Turkish by the TRMTEB group (BaysanGungor2025TR-MTEB). Table 11 reports the NDCG@10 scores achieved by each model. Across the TRMTEB retrieval benchmarks, TabiBERT consistently outperforms existing Turkish encoder models, achieving an overall NDCG@10 score of $75.44$ and surpassing prior monolingual models such as BERTurk. We attribute these gains in part to TabiBERT’s exposure to more diverse pre-training data, including bilingual sources, which appear to be particularly beneficial for retrieval-oriented tasks. TabiBERT attains the highest NDCG@10 scores on the BiText and MsMarco-TR datasets. On Scifact-TR, TabiBERT trails the best-performing model, YTU-Cosmos-BERT, by approximately $6$ points (best score: $79.88$). However, Scifact-TR is the smallest dataset in the TRMTEB suite, containing only $339$ test instances, and is therefore the least representative benchmark in the retrieval evaluation. For Fiqa-TR, TabiBERT achieves the second-highest NDCG@10 score of $50.67$, exceeding BERTurk by $1.21$ points. Similarly, on NFCorpus-TR, TabiBERT ranks second with an NDCG@10 score of $31.19$, only $1.28$ points below the best-performing model, BERTurk. On Quora-TR, all models except TurkishBERTweet, including mmBERT, exhibit nearly identical performance, with scores clustered around $92.50$, and TabiBERT performing slightly below this level.
Table 11: Downstream performance of models on Retrieval tasks. The number of test samples per task is also given. For each column, the highest score among the Turkish models (excluding mmBERT) is shown in bold.
Model BiText MsMarco-TR Scifact-TR Fiqa-TR NFCorpus-TR Quora-TR Weighted Avg
Test Size $3,000$ $31,692$ $339$ $1,706$ $12,334$ $15,675$
TurkishBERTweet $95.98$ $74.20$ $68.25$ $36.23$ $27.79$ $86.84$ $68.40$
YTU-Cosmos-BERT $97.25$ $81.99$ $79.88$ $51.10$ $28.36$ $92.87$ $74.29$
BERTurk $96.77$ $81.73$ $78.47$ $49.46$ $32.47$ $92.72$ $74.84$
TabiBERT $99.42$ $83.31$ $74.22$ $50.67$ $31.19$ $92.47$ $75.44$
mmBERT $99.49$ $83.68$ $80.00$ $52.29$ $33.44$ $92.78$ $76.20$
A.1.7 Code Retrieval
We evaluated code retrieval performance using four datasets: Apps-TR, CosQA-TR, StackOverflowQA-TR, CodeSearchNet-21K-subset-TR (TabiBERTCodeRetrievalDatasets). Table 12 presents the NDCG@10 scores obtained by each model. In code retrieval tasks, TabiBERT outperforms all other Turkish models, achieving an average NDCG@10 score of $56.95$, considerably higher than the next best result of $54.54$ by BERTurk. Thanks to its support for longer contexts and the incorporation of code in pre-training data, the model demonstrates clear strengths in code retrieval tasks. Across the individual tasks, TabiBERT secures the top NDCG@10 score in three out of four datasets, with being the second highest in the Apps-TR, lagging behind only by $0.69$ points.
Table 12: Downstream performance of models on Code Retrieval tasks. The number of test samples per task is also given. For each column, the highest score among the Turkish models (excluding mmBERT) is shown in bold.
Model Apps-TR CosQA-TR StackoverflowQA-TR CodeSearchNet-21K-TR Weighted Avg
Test Size $3,770$ $500$ $1,994$ $3,000$
TurkishBERTweet $04.48$ $82.32$ $67.01$ $70.41$ $43.49$
YTU-Cosmos-BERT $14.47$ $82.04$ $82.65$ $79.35$ $53.80$
BERTurk $18.18$ $85.72$ $79.93$ $78.15$ $54.54$
TabiBERT $17.49$ $88.11$ $82.85$ $84.12$ $56.95$
mmBERT $32.24$ $89.31$ $90.43$ $88.36$ $66.02$
A.1.8 Academic Domain Tasks
We evaluated academic domain tasks using four datasets: MedNLI-TR, PubmedRCT-10K-TR, SciCite-TR, ThesisAbstractClassification-11K (TabiBERTAcademicDomainDatasets). The MedNLI-TR dataset is an NLI dataset, the other three datasets are text classification datasets. Table 13 presents the results obtained by each model, reporting the macro F1 scores. We observe substantial improvements in academic domain tasks, such as NLI and multi-class text classification. For PubmedRCT-10K-TR task, TabiBERT shares the second place with YTU-Cosmos-BERT with a score of 75.32, only 0.29 behind the top performer BERTurk. On SciCite-TR, TabiBERT achieves a score of $83.29$, closely trailing YTU-Cosmos-BERT’s leading score of $84.09$ with a difference of $0.80$. On the demanding task of Thesis-Abstract-Classification-11K, characterized by long input contexts (with a 95th percentile context length of $1,152$ tokens) and sparse number of samples for each class (a sum of 60 examples per class throughout all splits), TabiBERT attains the highest score among Turkish models with $50.77$, outperforming the next best model BERTurk by $1.57$ points. Lastly, on the task of MedNLI-TR, TabiBERT achieves the highest score of $80.85$, slightly surpassing the second-best YTU-Cosmos-BERT by $0.23$ points. By leveraging a pre-training dataset rich in Turkish scientific articles and graduate theses, TabiBERT sets a new benchmark among Turkish models. These results highlight how TabiBERT’s exposure to academic domain texts and ability to process longer inputs underpin its strong performance in this area.
Table 13: Downstream performance of models on Academic domain tasks. The number of test samples per task is also given. For each column, the highest score among the Turkish models (excluding mmBERT) is shown in bold.
Model PubmedRCT-20K-TR SciCite-TR Thesis-Abstract-Classification-11K MedNLI-TR Weighted Avg
Task Type Text Clf Text Clf Text Clf NLI
Test Size $1,500$ $1,859$ $1,683$ $1,422$
TurkishBERTweet $70.07$ $79.59$ $31.18$ $72.05$ $63.12$
YTU-Cosmos-BERT $75.32$ $84.09$ $47.56$ $80.62$ $71.78$
BERTurk $75.61$ $81.60$ $49.20$ $79.90$ $71.40$
TabiBERT $75.32$ $83.29$ $50.77$ $80.85$ $72.44$
mmBERT $74.37$ $83.27$ $52.47$ $80.83$ $72.65$